how many grams of Kr are in a 5.69 L cylinder at 56.2 C and 3.91atm

Question 2 of 13> How many grams of Kr are in a 5.69 L cylinder at 56.2 °C and 3.91 atm?
Answer #1

Volume of the gas,V 5.69L Tempearture, T-56.2 °c =(56.2 +273)K = 329.2 K Gas consttnat(R)-0.0821L atm K mol1 Pressure , P-3.9